Fermi National Laboratory


Accelerator Update

Friday January 16
The day shift began with Operations monitoring store 3172, with MiniBooNE in access, with the Antiproton Source (Pbar) stacking, and with Recycler experts conducting studies.

At 9:34 AM, operators turned off Booster for an hour access. Operators turned the machine back on and resumed stacking at 11:01 AM.

Operations terminated store 3172 at 11:30 AM. Cryo system technicians began work on several wet engines around the Tevatron (TeV).

At 2:23 PM, Operations put the TeV into shot setup.

MiniBooNE finished their tunnel work and requested beam at 3:15 PM. An I- Source access, to inspect generator brushes, held off all beam until 4:18 PM.

At 4:23 PM, Operations established store 3175 with an initial luminosity of 47.08E30.

Operations setup the Recycler to receive shots of antiprotons at 4:33 PM.

Pbar resumed stacking at 5:27 PM.

At 9 PM, a Pbar Debuncher RF station (DRF1) and a kicker (D:IKIK) tripped off. Operators investigated and found a bad fuse on the RF station. An expert came in to check out the kicker. Stacking resumed at 10:55 PM.

Saturday January 17
The midnight shift began with Operations monitoring store 3175, with Pbar stacking, and with MiniBooNE taking beam.

At 6:29 AM, D:IKIK tripped off for the first of six times over the next hour. Operators asked an expert to investigate. He discovered a time-out problem and had it repaired by 9:38 AM.

A Linac RF station (KRF5) began tripping off at 3:57 PM. After following instructions from a Linac tech, operators got the station working and resumed stacking at 4:30 PM.

At 5:04 PM, Operations terminated store 3175.

By 5:24 PM, both the Collider Detector Facility (CDF) and D-zero (D0) were accessing their collision halls. They were both out of access by 7:27 PM.

Operations put the TeV into a dry squeeze at 7:55 PM and into shot setup at 10:11 PM.

At 11 PM, Operations established store 3177 with an initial luminosity of 48.58E30.

Sunday January 18
The midnight shift began with Operations monitoring store 3177, with MiniBooNE taking beam, and with Pbar stacking.

Main Injector (MI) BPMs caused problems at two locations around the MI ring at 1:33 AM. This will require further investigation during the day. (This problem was resolved by 1:33 PM by backing out of some new code that was installed by a system reboot.)

At 1:56 AM, Operations sent shots of antiprotons to the Recycler. Stacking resumed at 2:02 AM.

Operations finished a round of tuning to cleanup the MiniBooNE beam at 4:44 AM.

Operators reported that the cryo system cold compressor for TeV sector E2 was alarming on a power supply fault. The Cryo Coordinator will send someone out to investigate between stores.

MiniBooNE beam required another round of tuning at 10:13 AM.

Around 2 PM, a Switchyard (SY) expert conducted about a half hour worth of studies. A MI expert also began a study at this time. He finished his study around 7 PM.

At 8:28 PM, a SY expert established beam to the SY dump.

Operations terminated store 3177 at 9 PM. Its total integrated luminosity was 2020nb-1.

D0 experimenters accessed their collision hall 9:12 PM. They were out of access two hours later.

At 11:53 PM, Operations was having problems with the D0 low beta magnets. After a reboot and zeroing, they resumed normal functioning.

Monday January 19
The midnight shift began with Pbar stacking and with MiniBooNE taking beam.

At 12:31 AM, Operations put the TeV into shot setup, but had to terminate the proton store at 2:41 due to an RF problem; TRF7 was off. No antiprotons were in the machine.

Operations established store 3179 at 4:40 AM with an initial luminosity of 50.73E30.

At 5:32 AM, operators replaced a regulator for the I- Source due to excessive sparking. This helped a lot, but the source was still sparking.

Today's Plans
The plan for today is to continue with shots and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/20/2004   email Fermilab